Calculate the perimeter of the window


1
Expert's answer
2021-08-30T14:08:26-0400



The perimeter of the window above is equal to the perimeter of the rectangular part of the window + the perimeter of the semi circular part


perimeter of the window = rectangle perimeter + semi circle perimeter


perimeter of the rectangle = h + w + h + w = "2(h+w)"

perimeter of the semi circle = 1/2 × perimeter of circle = 1/2 × 2πr = "\pi r"


"\therefore" the perimeter of the window = "2(h+w) + \u03c0r"